OpenStudy (abhisar):

INCOMPLETE..PLS DON'T COMMENT A cocise note on Atomic Structure This tutorial will be focused on A-level numerical problems based on Atomic Structure

OpenStudy (abhisar):

\(\bf \large What~is~the~meaning~of~word~Atom?\) Atom is a greek word which means \(\bf Indivisible\) i.e an ultimate particle which cannot be further subdivided. During 1803-1808 \(\bf John~Dalton\) considered that all matter was composed of small particles called atoms. He gave some postulates about the structure of atom which we know as Dalton's Atomic theory.

OpenStudy (abhisar):

\(\sf \huge \color{green}{Dalton's~Atomic~theory}\) This theory is based on law of mass of conservation and law of definite proportion. The salient features of this theory are:- \(\bullet\) Each element is composed of extremely small particles called atoms. \(\bullet\) Atoms of particular element are like but differ from atom's of other element. \(\bullet\) Atom of each element is an ultimate particle and it has a characteristic mass but is structureless. \(\bullet\) Atoms are indestructible i.e they can neither be created nor be destroyed. \(\bullet\) Atoms take part in chemical reaction to form molecule.

OpenStudy (abhisar):

Now, towards the end of nineteenth century scientists performed various experiments and the observations established that `Atoms can be further subdivided into sub-atomic particles.` This concept was very different from the Dalton's atomic theory. Dalton was able to explain the law of conservation of mass, law of constant proportion very successfully however it failed to explain the results of other experiments, for example, it was known that substances like glass or ebonite when rubbed with silk or fur generate electricity. Many different kinds of sub-atomic particles like electron, proton, neutron were discovered in the twentieth century which also proved that Dalton's theory is wrong.

OpenStudy (abhisar):

In the coming years various Atomic models were proposed to explain the structure of atom and distribution of sub-atomic particles within it. Let us discuss few of them in brief \(\huge \sf \color{red}{Thomson~Model}\) \(\bf J.J~Thomson\), in 1898, proposed that an atom is a spherical structure in which the positive charge is uniformly distributed. The electrons are embedded into it in such a manner as to give the most stable electrostatic arrangement. This model was also called as \(\bf Water~Melon\) model due to the similarity in distribution. The melon part represents the positively charged sphere and the seeds represent the electrons.

OpenStudy (abhisar):

Thomson's model explained the neutrality of the atom but had following two drawbacks: \(\bullet\) It considered that mass of atom is uniformly distributed. This was proved wrong by Rutherford. \(\bullet\) It is a static model and does not reflect the movement of electron. \(\bullet\) According to this model, hydrogen can give rise to only one spectral line, contrary to the observed fact of several lines.

OpenStudy (abhisar):

\(\huge \sf \color{purple}{Rutherford's~\alpha -Scattering}\\ \huge \sf \color{purple}{Experiment}\) Now this is a topic that students ask me about frequently. It deals with three sub-topjcs. What was the experiment, what were the observations and what were the conclusions drawn. So let's check them one by one. In order to understand the actual experiment and the experimental setup watch \(\href{https:///www.youtube.com/watch?v=zUtIrO3fUgg}{\bf \underline{this~ video}}\) by Nick Duell. \(\sf \large \color{green}{Observations}\) \(\bullet\) Most of the \(\sf \alpha\)-particles (nearly 99.99%) went straight without suffering any deflections. \(\bullet\) Few \(\sf \alpha-particles\) got deflected through small angles. \(\bullet\) A very few \(\sf \alpha-particles\) (about one in 20, 000) did not pass through the foil at all but suffered large deflections (more than 90°) or even come back in the direction from which they came i.e a deflection of 180° \(\sf \large \color{red}{Conclusions}\) \(\bullet\) Since few of the \(\alpha\)-particles were defleted from their original path through moderate angles; it was cocluded that whole of the +ve charge is concentrated and the space occupied by this positive charge is very small in the atom. \(\Rightarrow\)Whenever \(\alpha\)-particle come closer to this point, they suffer a force of repulsion and deviate from their paths. \(\Rightarrow\)The positively charged heavy mass which occupies only a small volume in an atom is called \(\bf Nucleus\). It is supposed to be present at the centre of the atom. \(\bullet\) Since most of the \(\alpha\)-particle went straight through the metal foil undeflected, it means that there must be very large empty space within the atom. \(\bullet\) A very few of the \(\alpha\)-particles suffered strong deflections or even returned on their path indicating that the nucleus is rigid and \(\alpha\)-particles recoil due to direct collision with the heavy positively charged mass.

OpenStudy (abhisar):

Based on his observations and conclusions, he put forward an Atomic model whose main points are discussed below. \(\bf Rutherford's~Atomic~Model\) \(\bullet\) An atom has two parts, one is the centralised part called nucleus and the other is extra-nuclear part. \(\bullet\) Nucleus contains nucleons like positively charged protons. \(\bullet\) Radius of an atom is of order \(\sf 10^{-10}\)m and radius of Nucleus is of order \(\sf 10^{-15}\)m. This implies that radius of an atom is \(\sf 10^5\)times the radius of nucleus. \(\bf \color{red}{Drawbacks}\) \(\bullet\) According to \(\bf Maxwell's\) theory when a charged particle is celerated it loses its energy and hence it must be spiralled into the nucleus but practically it does not happen. \(\bullet\) Rutherford was not able to explain electronic configuration and spectrum.

OpenStudy (abhisar):

\(\bf \large\color{green}{\text{Bohr's Atomic Model for Hydrogen Atom}}\) \(\bf Neils~Bohr\) was the first to explain quantitatively the general features of hydrogen atom (or any other single electron species) and its spectrum. Bohr's model for hydrogen is based on the following postulates: \(\bullet\) The electron in the hydrogen atom can move around the nucleus in a circular path of fixed radius and energy. These paths are called as \(\bf Orbits\), stationary states or alowed energy states. These orbits are arranged concentrically around the nucleus. \(\bullet\) The energy of an electron in the orbit does not change with time. However, the electron will move from lower stationary sate to a higher stationary state when required amount of energy is absorbed by the electron or energy is emitted when electron moves from higher stationary state to lower. `The energy change does not take place in a continuous manner but occurs in discrete packets or` \(\bf Quantas\). \(\bullet\) Angular momentum of an electron in an orbit is integral multiple of \(\sf \huge \frac{h}{2\pi}\). \(\sf \Large ~~~~~~~~~~~~~~~~~~~~mvr=\huge \frac{nh}{2\pi}\), Where n is the number of orbit and h is `Plank's constant.` Thus an electron can move only in those orbits for which its angular momentum is integral multiple of h/2\(\sf \pi\)

OpenStudy (abhisar):

According to the Bohr's Atomic model for hydrogen atom: a) The stationary states for electron are numbered n=1,2,3.... These integral numbers are known as \(\bf Principal~Quantum~numbers\) b) The radius of any stationary state n in Angstroms unit can be expressed as, \(\sf \boxed{\Large r_n=\huge \frac{0.529\times n^2} {z}}\), where z is the atomic number of single electron species. 0.529 angstroms is the radius of the first stationary state of hydrogen atom and is called as \(\bf Bohr~Radius\). c) The energy associated with an electron in a particular stationary state n can be expressed as \(\sf \boxed{\Large E_n=-R_H(\frac{Z^2}{n^2})}\), where \(\sf R_H\) is called as Rydberg constant and its value is 2.18\(\sf \times 10^{-18}J\). d) Similarly, velocity of an electron associated with a stationary state n can be expressed as \(\sf \boxed{2.18\times 10^{6}\huge \frac{Z}{n}}m/s\)

OpenStudy (anonymous):

This is fascinating subject, I'm glad you did this, and it's very a good review, thanks Abhisar!

OpenStudy (abhisar):

I never got time to complete it....:D Thnx though :3